20. X received the following emoluments during the previous year ending March 31, 2021 basis salary 90,0000 dearness allowance 90000(net forming part of salary), bonus Re 74.000 and commission :40000
his employer provides the following perquisites:
A rent free furnished house in cochin: 41000
Lease rent of unfurnished house

Cot of steel furniture provided( written down value Rs.12,000, steel furniture is provided through the previous year) : 14000

Rent of furniture provided (rent in yet to be paid by employer): 17000 Motorcar expenditure of the employee (reimbursed by employer):6000
Determine the value of the perquisites in respect of rent-free furnished house if: (a)X is an officer of the Government of West Bengal and Rs. 18 000 is licence fee of the unfurnished house as per the State Government Rules, (b) X of the Indian Airlines, and (c) X is managing director of a private sector company .
1
Expert's answer
2021-10-12T16:20:52-0400

a)Where the accommodation is provided by the Central Government or any State Government to the employees either holding office or post in connection with the affairs of the Union or of such State, the value of perquisite in respect of such accommodation is equal to the licence fee, which would have been determined by the Central or State Government in accordance with the rules framed by the Government.

accommodation is equal to the licence fee+0,1cost of Furniture=18000+0.1(14000+17000+6000)=18000+3700=21700


b) Having population exceeding 25 lacs as per 2001 census:15% of salary for the period during which the employee occupied the said accommodation+0,1cost of Furniture

accommodation=0.15(90000+74000+40000)+0.1(14000+17000+6000)=34300


c)Having population exceeding 25 lacs as per 2001 census:15% of salary for the period during which the employee occupied the said accommodation+0,1cost of Furniture

accommodation=0.15(90000+74000+40000)+0.1(14000+17000+6000)=34300